Fermi National Laboratory


Accelerator Update for January 30, 2006 - February 1, 2006

Monday January 30

The day shift began with Operations monitoring the Antiproton Source (Pbar) stack and the Recycler stash, with Switchyard (SY) off due to an upcoming shot setup, and with NuMI and MiniBooNE taking beam.

At 8:23 AM, a cryo system expert reported that the Tevatron (TeV) sector C3 wet engine flywheel connecting rod had snapped and its piston was damaged. They began repairs. Operations dumped the TeV ramp for this work; no shot.

Operations prepared the Collider Detector at Fermilab (CDF) and D-Zero (D0) for collision hall accesses.

Operations began the TeV turn-on at 4:55 PM.

Pbar began stacking at 7:06 PM.

Operations put the TeV into shot setup at 8:27 PM.

Operations established store 4616 at 10:19 PM with an initial luminosity of 0.66E30. (This store was intentionally planned as a low luminosity store.)

Pbar resumed stacking at 11:58 PM.

Tuesday January 31

The midnight shift began with Operations monitoring store 4616 and the stash, and with Switchyard (SY), NuMI, and MiniBooNE taking beam.

Operations began preparing the Recycler to stash antiprotons at 5:23 AM. The stash ended and stacking resumed at 6:26 AM. The Meson MCenter experiment resumed taking beam.

The Recycler stashed antiprotons at 10:27 AM. The stash ended at 11:50 AM and stacking resumed.

Experts began working on the Tevatron Electron Lens (TEL) at 2:35 PM. The stash ended and stacking resumed at 5:47 PM. SY resumed taking beam.

The Recycler stashed antiprotons at 4:50 PM.

At 9:36 PM, a Linac Klystron (KRF1) tripped off on an overtemp. Operators investigated and changed an external fan.

Wednesday February 1

The midnight shift began with Operations monitoring store 4616 and the stash, with Pbar stacking, and with SY, NuMI, and MiniBooNE taking beam.

At 2:27 AM, the TEL tripped off for the first of many times during the rest of the shift. Operators reset it each time.

Operations terminated store 4616 at 5:28 AM.

At 5:36 AM, Operations reported that the Pbar Lithium Lens tripped off. Operators called in an expert to investigate.

Operations prepared CDF for access at 5:52 AM. They were out of access by 6:59 AM.

At 7:24 AM, the Lens power supply expert reported that he replaced a bad SCR and turned the supply back on, but there are more problems. Experts began discussing the situation.

The Plans for Wednesday

The plans are to investigate the Pbar Lens problem. Operations will establish a store with the stash this morning.

 
Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 02/01/2006    email Fermilab